OpenStudy (anonymous):

solve for theta between [0,2pi] sin(3theta)+sin(theta)=0

OpenStudy (zarkon):

one can show that \[\sin(3\theta)+\sin(\theta)=4\sin(\theta)\cos^2(\theta)\]